OpenStudy (anonymous):

Evaluate cot -220 degrees Round your answer to the nearest hundredth. A. -1.19 B. -0.84 C. 1.56 D. -1.31

OpenStudy (anonymous):

\[\cot \left( -220 \right)=-\cot 220=-\cot \left( 180+40 \right)=-\cot 40=-\cot \left( 90-50 \right)=-\tan 50\] =?

OpenStudy (anonymous):

-1.19

OpenStudy (anonymous):

\[-\cot \left( 90-50 \right)=-\tan 50 =?\]